John buys a new solar power energy grid worth $607,463 for his company's warehouse. The contract stipulates payments will use a nominal interest rate of 10.7% compounded monthly for 41 months. John agrees to pay $ 87 as down payment on the condition that the first payment will not be until the start of the 10 month if he pays interest on the current balance until he begins payments.
Determine the monthly payments.
Answer #1

Monthly payments
Case I=(607463-87)*(1+10.7%/12)^9*10.7%/12*1/(1-1/(1+10.7%/12)^41)=19228.24377

Case II=(607463-87)*(1+10.7%/12)^10*10.7%/12*1/(1-1/(1+10.7%/12)^41)=19399.69561

Case III=(607463-87)*(1+10.7%/12)^8*10.7%/12*1/(1-1/(1+10.7%/12)^41)=19058.3072